Jefferson Lab HALL A Logbook for March 05 - All EXCEPT Automatic Entries

  • Main INDEX

    Thursday

  • 142917 03/31/05 23:43 Ramesh T5 neutrons in new Kinematics
  • 142911 03/31/05 19:55 Ulmer Neutron rates
  • 142908 03/31/05 19:27 Ulmer Hall A Tools
  • 142907 03/31/05 19:25 Ulmer Neutron Rates
  • 142904 03/31/05 18:15 Ulmer Bigbite dE spectra
  • 142903 03/31/05 18:13 Ulmer Neutron Block ADCs
  • 142902 03/31/05 18:07 Ulmer Neutron Block TDCs
  • 142900 03/31/05 17:18 ran T5 protons TOF
  • 142890 03/31/05 16:05 Reitz Day Shift Summary
  • 142886 03/31/05 15:31 Moffit My access activity
  • 142878 03/31/05 09:16 ran N1 - L27 raised HV by 50 V
  • 142877 03/31/05 09:13 ran controled access -ROC 5 restarted
  • 142872 03/31/05 08:25 lkaufman Runs 3093-3095
  • 142869 03/31/05 08:06 lkaufman Owl Shift Summary
  • 142866 03/31/05 07:58 yijiang empty histrograms in online analysis gui for Bigbite
  • 142865 03/31/05 07:50 lkaufman Late Entry: N2 level low for right Q3, ok now
  • 142860 03/31/05 05:48 lkaufman Reboot TS2
  • 142857 03/31/05 04:42 lkaufman Back to 45 uA
  • 142852 03/31/05 03:37 lkaufman Hall A Tools Screen
  • 142851 03/31/05 03:35 lkaufman Neutron Detector Scalers
  • 142850 03/31/05 03:24 lkaufman Current from 45->40 uA
  • 142843 03/31/05 01:10 lkaufman N1 Scalers
  • 142841 03/31/05 00:47 lkaufman Q1 tripped and reset
  • 142839 03/31/05 00:08 mazouz adc spectrum for NA ADC 26
  • 142837 03/31/05 00:04 X. Zheng shift summary
  • Wednesday

  • 142836 03/30/05 23:57 mazouz few entries in one histogram
  • 142832 03/30/05 22:23 X. Zheng Q1 recovered
  • 142831 03/30/05 21:36 Reitz Q1 Screen after trip
  • 142830 03/30/05 21:31 Reitz Q1 Screen after trip
  • 142829 03/30/05 21:29 X. Zheng Left Q1 trip
  • 142823 03/30/05 20:00 Reitz BB/Neutron Scalers
  • 142822 03/30/05 19:58 X. Zheng Compton rate screen
  • 142821 03/30/05 19:57 X. Zheng Hall A tools screen
  • 142817 03/30/05 18:37 X. Zheng vacuum problem, reduce beam current to 45uA
  • 142809 03/30/05 16:04 LeRose Day Shift Summary
  • 142807 03/30/05 15:55 Adaq Target snapshot
  • 142802 03/30/05 14:59 suleiman lumi study for qweak
  • 142801 03/30/05 14:52 LeRose Hall A Tools
  • 142799 03/30/05 14:24 LeRose Back in business
  • 142798 03/30/05 13:57 LeRose Controlled Access
  • 142797 03/30/05 13:35 Ramesh Reason for peak at 400 channel
  • 142796 03/30/05 12:50 LeRose SLI
  • 142795 03/30/05 12:48 ran T5 neutron tof
  • 142792 03/30/05 12:25 adaq BB/ND scalers at 50uA
  • 142791 03/30/05 12:02 LeRose LUMI studies
  • 142790 03/30/05 11:02 LeRose Controlled Access
  • 142789 03/30/05 10:38 LeRose Right Magnets Failure
  • 142787 03/30/05 10:07 LeRose Beam Back
  • 142786 03/30/05 10:07 ran current increased to 50 microAmps
  • 142784 03/30/05 09:41 LeRose Compton
  • 142783 03/30/05 09:02 LeRose Beam Off
  • 142781 03/30/05 08:22 ran singles/coinc time issue
  • 142778 03/30/05 08:02 lkaufman Owl shift summary
  • 142774 03/30/05 06:39 ran pmiss distribution inside and outside ptof peak
  • 142772 03/30/05 06:17 ran command line used for last ptof
  • 142770 03/30/05 05:49 ran high Pmiss coinc protons
  • 142761 03/30/05 03:50 ran halla A tool screen shots
  • 142760 03/30/05 03:47 ran ND & BB scalers
  • 142759 03/30/05 03:46 ran target screens
  • Tuesday

  • 142748 03/29/05 23:56 Punjabi Shift Summary
  • 142738 03/29/05 20:21 Punjabi detector scalar screen
  • 142737 03/29/05 20:13 Punjabi Hall A tools
  • 142725 03/29/05 16:05 Higinbotham Check E Hit #1
  • 142723 03/29/05 16:00 Higinbotham Shift Summary
  • 142716 03/29/05 13:37 Higinbotham Hall A General Tools Screen
  • 142715 03/29/05 13:35 Higinbotham BB and ND Scalers
  • 142711 03/29/05 12:10 Higinbotham BeO and 12C Checks
  • 142710 03/29/05 12:07 Nilanga scaler rates with slanted 12C
  • 142709 03/29/05 11:58 NIlanga scaler rates
  • 142690 03/29/05 08:00 Ramesh Subed Shift Summary
  • 142684 03/29/05 06:13 ran beam is back
  • 142683 03/29/05 06:05 ran RF problems
  • 142679 03/29/05 04:56 ran compton DAQ problems
  • 142678 03/29/05 04:47 ran target screens
  • 142675 03/29/05 04:29 Ramesh/Ran T5 neutrons (high vs low pmiss)
  • 142674 03/29/05 04:29 Ramesh/Ran protons (high vs low pmiss)
  • 142667 03/29/05 01:16 ran protons TOF
  • 142664 03/29/05 00:55 ran T5 protons tof
  • 142663 03/29/05 00:50 Adaq Hall A tools
  • 142662 03/29/05 00:29 Adaq BB/ND scalers
  • 142658 03/29/05 00:03 Nilanga Shift Summary
  • Monday

  • 142647 03/28/05 18:16 X. Zheng plot for NA and BB screen
  • 142646 03/28/05 18:15 X. Zheng NA and BB screen
  • 142645 03/28/05 18:13 X. Zheng Hall A tools screen
  • 142644 03/28/05 18:11 X. Zheng Compton chart
  • 142641 03/28/05 15:57 Ole Hansen Shift Summary
  • 142640 03/28/05 13:32 R. Feuerbach Shifting of coincidence time
  • 142639 03/28/05 13:22 Ulmer Hall A Tools
  • 142638 03/28/05 12:53 Ulmer Going to beam permit
  • 142637 03/28/05 12:22 Ulmer Q1-Left trip
  • 142636 03/28/05 08:47 hyao status of target
  • 142634 03/28/05 08:08 ran target screens
  • 142633 03/28/05 08:07 ran target screen
  • 142631 03/28/05 08:01 Ramesh Shift Summary
  • 142628 03/28/05 07:27 ran corrected triple of the day
  • 142625 03/28/05 06:54 ran triple of the day
  • 142617 03/28/05 02:44 ran TS0 & ROC1 reboot
  • 142616 03/28/05 02:41 ran Q1 trip
  • 142613 03/28/05 01:27 Adaq Hall A tools
  • 142612 03/28/05 01:11 Adaq BB/ND scaler
  • Sunday

  • 142608 03/27/05 23:59 sawatzky Shift/Run Summary
  • 142599 03/27/05 20:51 sawatzky General Tools Screen
  • 142596 03/27/05 20:49 sawatzky Compton rates
  • 142595 03/27/05 20:43 Wang neutron adc /tdc Run 3003
  • 142594 03/27/05 20:22 sawatzky Target screens
  • 142593 03/27/05 20:20 sawatzky BB and NA rate screen
  • 142592 03/27/05 20:10 sawatzky Power cycled BB VME crate
  • 142591 03/27/05 20:09 sawatzky LHRS He level dropped below 70% briefly
  • 142587 03/27/05 19:42 sawatzky BB QCD initialization error
  • 142580 03/27/05 19:11 sawatzky Access to reset Neutron arm HV crate
  • 142579 03/27/05 18:40 sawatzky Neutron HVs tripped
  • 142577 03/27/05 18:30 sawatzky Poor beam, notified MCC
  • 142572 03/27/05 16:35 sawatzky Compton stripchart rates flat-lined
  • 142569 03/27/05 16:08 johna Shift Summary
  • 142565 03/27/05 14:49 hyao beam is back
  • 142563 03/27/05 14:19 hyao Beam is off
  • 142557 03/27/05 12:48 Ibrahim General Tools Screen
  • 142556 03/27/05 12:44 Ibrahim Neutron Detector and BigBite Single Rates
  • 142555 03/27/05 12:36 Ibrahim Compton Stripchart
  • 142550 03/27/05 10:48 Ramesh T5 neutrons in new kinematics
  • 142547 03/27/05 09:49 Ramesh Coincidence time changed in two run periods
  • 142541 03/27/05 08:14 hyao status of target
  • 142540 03/27/05 08:12 ramesh owl shift summary
  • 142538 03/27/05 07:26 ran T5 proton plots
  • 142537 03/27/05 07:24 ran T5 protons
  • 142528 03/27/05 04:12 adaq BB/ND scaler
  • 142527 03/27/05 03:36 Adaq hall A tools
  • 142526 03/27/05 03:27 Ramesh/Ran Plots of angles for two kinematics
  • Saturday

  • 142516 03/26/05 23:54 johna Shift summary
  • 142508 03/26/05 21:29 johna Automatic BTA software
  • 142506 03/26/05 21:24 ran day shift summary [late submission]
  • 142501 03/26/05 19:38 johna Hall A General Tools (snapshot)
  • 142500 03/26/05 19:37 johna scaler rates (snapshot)
  • 142491 03/26/05 17:01 reinhold minor target alarm
  • 142490 03/26/05 16:14 reinhold target status
  • 142485 03/26/05 15:23 ran halla tool screen snap-shot
  • 142484 03/26/05 15:18 ran scaler rates
  • 142465 03/26/05 10:48 R. Michaels Efficiency of R-arm too
  • 142464 03/26/05 10:31 R. Michaels Left arm scint. efficiency
  • 142463 03/26/05 10:13 hyao Status of target
  • 142462 03/26/05 09:43 ran ROC 5 & ROC 1 rebooted
  • 142458 03/26/05 08:13 P. Monaghan Owl shift summary
  • 142456 03/26/05 08:05 P. Monaghan Compton stripchart better
  • 142453 03/26/05 07:36 P. Monaghan Hall A controls screen
  • 142452 03/26/05 07:34 P. Monaghan Target charts
  • 142451 03/26/05 07:32 P. Monaghan BigBite/Neutron detector scalers
  • 142448 03/26/05 06:48 P. Monaghan Compton not looking good?
  • 142440 03/26/05 03:27 P. Monaghan Neutron detector HV crate reset
  • 142438 03/26/05 02:46 P. Monaghan Lost compton stripchart
  • 142435 03/26/05 02:07 P. Monaghan BigBite/Neutron detector scalers
  • 142434 03/26/05 01:58 P. Monaghan Compton frozen?
  • Friday

  • 142426 03/25/05 23:05 johna Compton rates look good (snapshot)
  • 142421 03/25/05 21:34 johna (minor) problems restarting Compton run
  • 142414 03/25/05 18:58 johna Hall A General Tools (snapshot)
  • 142413 03/25/05 18:57 johna Neutron detector single rates (snapshot)
  • 142404 03/25/05 16:15 reinhold target status
  • 142403 03/25/05 16:11 saw Kinematics
  • 142401 03/25/05 16:02 Fatiha Shift Summary
  • 142399 03/25/05 15:15 Fatiha Shift Summary
  • 142393 03/25/05 13:23 Ulmer Neutron ADCs
  • 142392 03/25/05 12:50 Nanda Compton Rates
  • 142390 03/25/05 12:30 Ulmer Neutron ADC shows change
  • 142388 03/25/05 11:56 Ulmer Left Arm HV
  • 142387 03/25/05 11:56 Ulmer Right Arm HV
  • 142386 03/25/05 11:54 Fatiha BigBite single rates
  • 142385 03/25/05 11:54 Ulmer Neutron HV
  • 142384 03/25/05 11:53 Ulmer Bigbite HV
  • 142383 03/25/05 11:52 Ulmer Hall A Tools
  • 142382 03/25/05 11:47 Fatiha Magnet Screen
  • 142381 03/25/05 11:35 Fatiha ps1 changes
  • 142376 03/25/05 11:21 Ulmer Harp continued
  • 142375 03/25/05 11:20 Ulmer Harp Scan
  • 142373 03/25/05 08:50 Ramesh T5 neutrons in new kinematics
  • 142372 03/25/05 08:42 Ulmer PR Layer 2 Block 13
  • 142370 03/25/05 08:25 Fatiha Beam studies
  • 142368 03/25/05 07:54 markowitz tools screen
  • 142367 03/25/05 07:53 markowitz neutron scalers screen
  • 142354 03/25/05 03:39 saw ND/BB Scaler rates for 40uA
  • 142346 03/25/05 00:36 saw Upper neutron HV crate
  • 142344 03/25/05 00:27 Adaq HV problems - run 2916
  • 142343 03/25/05 00:26 johna Shift summary
  • Thursday

  • 142337 03/24/05 22:07 johna Compton rates
  • 142336 03/24/05 22:06 johna Hall A General Tools
  • 142335 03/24/05 22:05 johna Neutron detector single rates: snapshot
  • 142334 03/24/05 21:46 johna Harp scan
  • 142332 03/24/05 18:34 johna status = no beam
  • 142331 03/24/05 16:40 Ramesh/Larry T5 neutrons with new code
  • 142329 03/24/05 16:01 weinstein shift summary
  • 142327 03/24/05 15:34 Sirish Beam Halo
  • 142326 03/24/05 15:13 X. Jiang TS1 hang. Reboot
  • 142324 03/24/05 15:11 Olivier Triple coincidence neutron spectrum
  • 142319 03/24/05 13:30 potokar Hall A General Tools Snapshot
  • 142318 03/24/05 13:28 potokar Neuron Detector Single Rates Snapshot
  • 142314 03/24/05 12:24 Sirish Compton Tune
  • 142310 03/24/05 10:05 potokar Hall A General Tools Snapshot
  • 142309 03/24/05 10:01 X. Jiang HV screen. Neutron lower crate.
  • 142308 03/24/05 09:59 X. Jiang Scaler creen at run 2907
  • 142302 03/24/05 08:42 saw QDC id 0 errrors fixed
  • 142298 03/24/05 08:06 saw QDC id 0 not initialized errors are bad
  • 142294 03/24/05 07:56 Markowitz shift summary
  • 142291 03/24/05 07:52 saw OTR was in
  • 142290 03/24/05 07:46 saw QDC id 0 not initialized errors
  • 142285 03/24/05 07:23 markowitz hall a tools screen grab
  • 142284 03/24/05 07:22 markowitz neutron rates at 15 uA
  • 142283 03/24/05 07:11 saw Neutron detector and Bigbite rates
  • 142282 03/24/05 07:08 saw Rebooted ROC5
  • 142280 03/24/05 06:44 markowitz wrong angle in last entry
  • 142278 03/24/05 06:32 saw Singles rates, beam quality
  • 142275 03/24/05 06:14 markowitz change kinematics to 36.8, 1.42 GeV
  • 142274 03/24/05 05:36 saw Change target to slanted carbon
  • 142273 03/24/05 05:23 markowitz back out of compton again
  • 142272 03/24/05 03:50 Markowitz mcc asked for compton back
  • 142271 03/24/05 02:47 Markowitz MCC gives up on compton
  • Wednesday

  • 142268 03/23/05 23:56 Sirca SWING shift summary
  • 142267 03/23/05 23:42 Sirca Compton instructions for night shift
  • 142266 03/23/05 22:04 Nanda Compton Rates
  • 142265 03/23/05 21:55 Sirca Beam being retuned in injector and linac
  • 142264 03/23/05 18:30 folts tech
  • 142263 03/23/05 18:29 johna Work on Q2 finished
  • 142261 03/23/05 16:57 Olivier More notes about the new physics module
  • 142260 03/23/05 16:54 Olivier Use of new physics module
  • 142259 03/23/05 16:41 Olivier S/N improvement for deuterium x>1
  • 142257 03/23/05 16:00 LeRose Day Shift Summary
  • 142256 03/23/05 14:46 LeRose Hall A Tools
  • 142255 03/23/05 14:45 LeRose Neutron Detector Scalers
  • 142254 03/23/05 13:27 LeRose SCR defined
  • 142253 03/23/05 12:44 LeRose TLA confusion
  • 142252 03/23/05 12:22 LeRose Left Q2 Update
  • 142251 03/23/05 11:58 Olivier SK.emiss or SK.tb ?
  • 142250 03/23/05 11:52 LeRose Left Q2 Still Down
  • 142249 03/23/05 10:35 ran T5 protons & neutrons TOF .
  • 142248 03/23/05 08:56 LeRose Left Q2 Still Down
  • 142247 03/23/05 08:06 Markowitz shift summary
  • 142245 03/23/05 07:00 Markowitz Q2 woes continue, awaiting Tech help
  • 142242 03/23/05 04:36 saw Q2 tripped again
  • 142238 03/23/05 03:47 Folts tech
  • 142237 03/23/05 03:35 Markowitz tech called in for Q2
  • 142235 03/23/05 01:11 saw More Q2 trips
  • 142234 03/23/05 00:34 Markowitz Q2 reset
  • 142232 03/23/05 00:06 Sirca SWING shift summary
  • 142231 03/23/05 00:02 Sirca Hall access to reset HRSL Q2
  • Tuesday

  • 142230 03/22/05 23:53 Ramesh Tripple coinc. neutrons
  • 142229 03/22/05 23:39 Sirca HRS Left Q2 trip
  • 142219 03/22/05 20:24 eip emiss limits
  • 142218 03/22/05 20:16 Sirca Neutron det single rates snapshot
  • 142217 03/22/05 20:15 Sirca Hall A General Tools snapshot OK
  • 142216 03/22/05 20:12 Sirca Hall A General Tools snapshot
  • 142215 03/22/05 20:09 Sirca Resume DAQ
  • 142213 03/22/05 19:36 Sirca HRS Left Q1 restored
  • 142212 03/22/05 18:56 Sirca Hall access to reset HRS Left Q1
  • 142211 03/22/05 18:48 Sirca Tech: HRS Left Q1 trip
  • 142207 03/22/05 17:40 Sirca HRS Left Q2 trips again
  • 142203 03/22/05 16:45 Sirca HRS Left Q2 restored
  • 142201 03/22/05 16:32 weinstein neutron cuts from (e,e'p) Emiss
  • 142200 03/22/05 16:29 Sirca HRS Left Q2 trip, strange phenomena during ramp-up
  • 142198 03/22/05 16:18 Sirca Neutron det single rates snapshot
  • 142196 03/22/05 16:07 weinstein shift summary - day
  • 142193 03/22/05 15:39 potokar Hall A General Tools Snapshot
  • 142180 03/22/05 11:16 R. Michaels new stage disks, rebooted adaql2
  • 142177 03/22/05 10:38 potokar Hall A General Tools Snapshot
  • 142176 03/22/05 10:31 R. Michaels Neutron HV check
  • 142175 03/22/05 10:27 potokar Neutron Det. Single Rates Snapshot
  • 142174 03/22/05 10:24 weinstein beam current back to 40uA
  • 142173 03/22/05 10:14 R. Michaels scaler code updated, history ntuple added
  • 142168 03/22/05 08:48 weinstein intermittent Hall A alarm
  • 142167 03/22/05 08:44 weinstein HRS-R angle not recorded
  • 142163 03/22/05 08:07 Markowitz Owl shift summary
  • 142160 03/22/05 07:09 markowitz neutron screens at 35 uA
  • 142159 03/22/05 07:08 markowitz bigbite striptool screen
  • 142158 03/22/05 07:08 Markowitz tools screen
  • 142142 03/22/05 00:00 Sirca Swing shift summary
  • Monday

  • 142139 03/21/05 23:11 Sirca I_beam 20 uA to 30 uA
  • 142138 03/21/05 22:45 Sirca Acc vacuum problem, reduce I_beam
  • 142133 03/21/05 21:31 Ramesh T5 neutrons
  • 142130 03/21/05 20:45 Sirca (Small) BB field fluctuation
  • 142129 03/21/05 20:38 Sirca Neutron det single rates snapshot
  • 142128 03/21/05 20:37 Sirca General Tools, HRSR angle readout
  • 142127 03/21/05 20:33 Sirca Resume production, BPM readouts
  • 142125 03/21/05 20:24 Sirca Harp scan results
  • 142124 03/21/05 19:00 Sirca Hall back to beam permit
  • 142123 03/21/05 18:29 Sirca Hall access to power-cycle HV
  • 142115 03/21/05 16:04 jpchen Day shift summary
  • 142114 03/21/05 15:39 jpchen asked MCC to ahve beam not through Compton Chicane
  • 142110 03/21/05 15:24 potokar Hall A General Tools Snapshot
  • 142109 03/21/05 14:47 jpchen run 2850-2852 wrong prescale factors
  • 142101 03/21/05 14:37 X. Jiang scaler screen with 40 uA beam
  • 142100 03/21/05 14:36 jpchen right dipole ready
  • 142099 03/21/05 14:27 X. Jiang Big Bite scaler screen
  • 142098 03/21/05 14:26 X. Jiang BigBite scaler screen with slant c12 target, 5 uA beam.
  • 142097 03/21/05 13:32 jpchen Harp scan
  • 142096 03/21/05 13:26 jpchen beam back. 5uA first for Harp scan
  • 142092 03/21/05 10:05 ran neutron tof
  • 142091 03/21/05 09:42 ran R HRS phi vs P miss with a cut on Mrec
  • 142090 03/21/05 09:16 jpchen Hall in controlled access
  • 142088 03/21/05 08:49 X. Jiang Take down all magnets.
  • 142087 03/21/05 08:49 jpchen beam study for 6 hours
  • 142086 03/21/05 08:47 X. Jiang Big Bite magnet control screen at the end of run 2847.
  • 142085 03/21/05 08:44 X. Jiang Magnet screen at end of run #2847
  • 142082 03/21/05 08:11 jpchen vacuum pump tripped, minimum 30 minutes no beam
  • 142079 03/21/05 08:05 Ramesh Shift Summary
  • 142076 03/21/05 06:43 ran R HRS phi vs P miss with a cut on Mrec
  • 142071 03/21/05 05:32 ran R gold phi vs P.miss phi
  • 142069 03/21/05 04:55 ran R arm in-plane-angle acceptance vs kinematics
  • 142063 03/21/05 02:26 Ramesh TS2 (SRC trig. supervisor) failed, rebooted, now OK
  • 142059 03/21/05 01:38 Adaq Target Status
  • 142058 03/21/05 01:32 adaq Hall A tools
  • 142057 03/21/05 01:31 adaq BB/ND scaler
  • Sunday

  • 142051 03/20/05 23:58 X. Zheng Swing shift summary
  • 142042 03/20/05 20:26 X. Zheng Running on Hall C laser now
  • 142039 03/20/05 19:47 X. Zheng beam back
  • 142037 03/20/05 19:34 X. Zheng source problem, lost beam
  • 142035 03/20/05 19:26 Sirca Neutron det single rates snapshot
  • 142034 03/20/05 19:24 Sirca Hall A General Tools snapshot
  • 142032 03/20/05 18:56 Olivier 192 subnet switch power cycled
  • 142027 03/20/05 17:55 X. Zheng teck - Left Q3 cryo unstable?
  • 142026 03/20/05 17:46 X. Zheng SRC ROC5&6 problem, controlled access
  • 142025 03/20/05 16:01 reinhold Shift summary
  • 142021 03/20/05 14:59 reinhold target alarm
  • 142018 03/20/05 14:04 potokar Neutron Detector Single Rates Snapshot
  • 142017 03/20/05 14:00 potokar HallA General Tools Snapshot
  • 142014 03/20/05 12:54 potokar Strip Tool Graph Window Snapshot
  • 142012 03/20/05 12:16 folts tech
  • 142009 03/20/05 10:56 reinhold Beam back at 40 uA
  • 142008 03/20/05 10:53 reinhold MCC calls: will send beam
  • 142007 03/20/05 10:48 reinhold Hall back to BEAM PERMIT
  • 142006 03/20/05 10:38 folts tech
  • 142005 03/20/05 10:24 reinhold Hall goes to power permit
  • 142004 03/20/05 09:39 reinhold All HRS left magnets set again
  • 142003 03/20/05 09:20 reinhold Q1 ramping up
  • 142002 03/20/05 08:39 reinhold Q1 soft reset fails
  • 142001 03/20/05 08:38 reinhold waiting for access and Ed
  • 142000 03/20/05 08:24 reinhold Q1 crash/trip
  • 141998 03/20/05 08:22 reinhold target status OK
  • 141997 03/20/05 08:08 Ramesh Shift Summary
  • 141991 03/20/05 06:53 Ramesh Analysis Computer terminals frozen
  • 141990 03/20/05 06:15 Ramesh Q1 back after mannual reset in escorted access
  • 141988 03/20/05 05:41 Ramesh Escorted access to reset Q1 (third time)
  • 141987 03/20/05 05:31 Ramesh Techs - Q1 again crashed
  • 141985 03/20/05 05:26 Ramesh Techs
  • 141982 03/20/05 04:53 adaq Target snapshot
  • 141981 03/20/05 04:39 adaq BB/ND scalers
  • 141980 03/20/05 04:38 adaq BB/ND scalers
  • 141979 03/20/05 04:37 adaq Hall A tools
  • 141975 03/20/05 04:32 Ramesh Q1 back after mannual rese
  • 141974 03/20/05 04:18 Ramesh Q1 crashed again and needed escorted access
  • 141969 03/20/05 02:40 Ramesh Q1 back
  • 141968 03/20/05 01:29 Ramesh controlled access to reset Q1
  • 141967 03/20/05 00:51 Ramesh left Q1 tripped
  • 141962 03/20/05 00:00 Sirca SWING shift summary
  • Saturday

  • 141961 03/19/05 23:41 Sirca Neutron det single rates snapshot
  • 141960 03/19/05 23:39 Sirca Hall A General Tools snapshot
  • 141955 03/19/05 22:25 Sirca Beam off for another 10-15 min
  • 141953 03/19/05 21:34 Sirca Beam off for injector work
  • 141945 03/19/05 19:38 Sirca Neutron det single rates snapshot
  • 141944 03/19/05 19:37 Sirca Hall A General Tools snapshot
  • 141937 03/19/05 16:38 hyao beam is back
  • 141934 03/19/05 16:32 hyao target status at 16:31
  • 141933 03/19/05 16:05 Ole Hansen Shift Summary
  • 141930 03/19/05 16:01 potokar Hall A General Tools Snapshot
  • 141924 03/19/05 14:15 potokar Neutron detector Single Rates Snapshot
  • 141923 03/19/05 14:11 Potokar Hall A General Tools Snapshot
  • 141918 03/19/05 12:33 Ole Hansen 40 uA restored
  • 141917 03/19/05 12:19 Ole Hansen BB scaler screen
  • 141916 03/19/05 12:19 Ole Hansen Hall A tools
  • 141910 03/19/05 08:17 reinhold Target status
  • 141905 03/19/05 07:58 Ramesh Shift Summary
  • 141904 03/19/05 07:51 Kolarkar Target Screens - Owl Shift
  • 141893 03/19/05 04:21 adaq Hall A tools
  • 141892 03/19/05 04:21 adaq BB scaler display
  • 141891 03/19/05 04:19 Ramesh 20 uA beam using Hall C laser
  • 141881 03/19/05 00:03 Sirca SWING shift summary
  • Friday

  • 141874 03/18/05 22:22 Sirca Neutron det single rates snapshot
  • 141873 03/18/05 22:21 Sirca Hall A General Tools snapshot
  • 141865 03/18/05 20:28 Sirca Key to 3rd arm electronics cage
  • 141860 03/18/05 18:58 Sirca Neutron det single rates snapshot
  • 141859 03/18/05 18:57 Sirca Hall A General Tools snapshot
  • 141858 03/18/05 18:56 Sirca Tg flammable gas sensors
  • 141857 03/18/05 18:45 Sirca Power glitch (?)
  • 141851 03/18/05 16:36 hyao swing shift: target status
  • 141849 03/18/05 16:19 Sirca Neutron det single rates snapshot
  • 141848 03/18/05 16:18 Sirca Hall A General Tools snapshot
  • 141846 03/18/05 16:11 Fatiha ABU screen
  • 141845 03/18/05 16:02 Fatiha Shift Summary
  • 141836 03/18/05 12:58 Fatiha Main control screen
  • 141835 03/18/05 12:57 Fatiha Bigbite single rates screen
  • 141832 03/18/05 11:58 R. Michaels auto ABU working again (but check it)
  • 141831 03/18/05 11:46 Fatiha Beam permit
  • 141829 03/18/05 11:39 saw Aux threshold lowered
  • 141828 03/18/05 11:13 Fatiha Controlled access
  • 141826 03/18/05 10:54 gilman cryotarget status deuterium loop
  • 141825 03/18/05 10:41 Fatiha TS2 restarted
  • 141818 03/18/05 09:14 Fatiha SRC HV reset
  • 141817 03/18/05 09:02 yijiang distribution of "Y1" on p2-pos(1-12)
  • 141814 03/18/05 08:08 Ramesh Subed Shift Summary
  • 141809 03/18/05 07:27 Ramesh HV GUI of ND/BB crashed
  • 141808 03/18/05 07:25 Kolarkar Target Screens - Owl shift
  • 141805 03/18/05 06:29 Ramesh BB tdc spectra ok
  • 141800 03/18/05 04:55 Adaq BB tdc spectra (all plane low)
  • 141799 03/18/05 04:45 Ramesh low tdc values in BB
  • 141791 03/18/05 02:43 Ramesh Deadtime fluctuating ramdomly
  • 141788 03/18/05 01:48 potokar netron detector single rates snapshot
  • 141787 03/18/05 01:45 potokar Hall A General Tools Snapshot
  • 141786 03/18/05 01:42 potokar Hall A General Tools Snapshoy
  • 141780 03/18/05 00:04 Sirca/X. Zhe HV trip during 2751?
  • 141779 03/18/05 00:01 X. Zheng swing shift summary
  • 141778 03/18/05 00:00 X. Zheng run #2752 had 190K events
  • Thursday

  • 141777 03/17/05 23:43 X. Zheng no beam for 15 minutes
  • 141775 03/17/05 23:37 Sirca/X. Zhe Neutron detector singles frozen/HV trip
  • 141774 03/17/05 23:37 Sirca Neutron det snapshot after HV reset
  • 141771 03/17/05 23:12 Sirca Hall A General Tools snapshot
  • 141754 03/17/05 16:39 Sirca Neutron det scalers snapshot
  • 141753 03/17/05 16:36 Sirca Hall A general tools snapshot
  • 141752 03/17/05 16:24 reinhold Target status OK
  • 141750 03/17/05 16:00 LeRose Day Shift Summary
  • 141749 03/17/05 15:21 slifer ND/BB scalers
  • 141747 03/17/05 15:13 LeRose Harp Scan
  • 141746 03/17/05 15:02 slifer target screens
  • 141745 03/17/05 15:00 P. Monaghan Changes to HV on neutron detector
  • 141744 03/17/05 14:59 LeRose Tools Screen/Right Dipole
  • 141743 03/17/05 14:57 slifer target to slanted C
  • 141741 03/17/05 14:17 slifer MCC update
  • 141740 03/17/05 13:11 slifer target to empty position
  • 141739 03/17/05 12:46 ran recoil mass
  • 141738 03/17/05 12:42 ran T5 coinc neutrons TOF in the neutron array
  • 141737 03/17/05 12:36 ran BigBite triple coinc protons
  • 141735 03/17/05 10:54 P. Monaghan Swapped pmts bases on neutron detector
  • 141733 03/17/05 10:29 LeRose Restricted Access
  • 141730 03/17/05 07:59 B. Craver Owl Shift Summary
  • 141729 03/17/05 07:56 B. Craver beam studies
  • 141724 03/17/05 06:25 Kolarkar Target Screens
  • 141723 03/17/05 06:23 Adaq General Tools screen shot
  • 141722 03/17/05 06:21 B. Craver neutron detector screen shot
  • 141717 03/17/05 04:15 Kolarkar Neutron Detector Single Rates Screen
  • 141710 03/17/05 01:30 eip detectrs replay
  • 141704 03/17/05 00:00 jpchen shift summary
  • Wednesday

  • 141701 03/16/05 23:39 V. Sulkosky ND Scalers
  • 141698 03/16/05 22:18 V. Sulkosky Hall A General Tools
  • 141697 03/16/05 22:16 jpchen TS1 not responding during run 2729
  • 141693 03/16/05 20:35 V. Sulkosky Neutron Detector Scalers
  • 141690 03/16/05 19:40 jpchen cryotarget status(3)
  • 141689 03/16/05 19:39 jpchen cryotarget status(2)
  • 141688 03/16/05 19:38 jpchen cryotarget status (1)
  • 141685 03/16/05 19:07 yijiang the strange distr. of preshower ADCs 28&35 still exist in run 2725
  • 141681 03/16/05 18:17 V. Sulkosky Neutron Detector Scalers
  • 141680 03/16/05 17:57 Yijiang strange distr. on preshower ADC 28&35
  • 141678 03/16/05 16:51 jpchen no beam, MCC run LEM procedure and then vacuum fault
  • 141676 03/16/05 16:33 jpchen beam tuning for Compton during run 2723
  • 141674 03/16/05 16:02 Boeglin Shift Summary
  • 141671 03/16/05 15:35 R. Michaels spare TDCs put in cabinet
  • 141670 03/16/05 15:18 Boeglin Hall A general Tools screen
  • 141669 03/16/05 15:17 Boeglin n-det scaer screen
  • 141667 03/16/05 15:11 R. Michaels T4 rates
  • 141666 03/16/05 15:00 R. Michaels Scaler counts vs time.
  • 141665 03/16/05 14:45 Boeglin Q1L trip
  • 141647 03/16/05 09:00 sawatzky Change in neutron bar tdc spectra
  • 141641 03/16/05 08:01 B. Craver Owl Shift Summary
  • 141640 03/16/05 07:45 ran coinc protons from C12
  • 141635 03/16/05 05:08 Adaq Target status
  • 141634 03/16/05 05:04 B. Craver General Tools screen shot
  • 141633 03/16/05 05:03 B. Craver neutron detector screen shot
  • 141624 03/16/05 01:23 Ramesh T6 rate (-)ve
  • Tuesday

  • 141618 03/15/05 23:55 perdrisat journal, swing shift
  • 141615 03/15/05 22:36 perdrisat fast feedback is off
  • 141612 03/15/05 22:20 perdrisa neutron single rate
  • 141605 03/15/05 19:24 mazouz Hall A General Tools
  • 141602 03/15/05 18:29 mazouz neutron detector single rates
  • 141599 03/15/05 17:36 R. Michaels problem of lacking spares
  • 141597 03/15/05 17:04 R. Feuerbach S2m L-HRS fixed itself?
  • 141590 03/15/05 16:02 boeglin Shift Summary
  • 141588 03/15/05 15:26 R. Michaels problems -- S2m, etc
  • 141587 03/15/05 15:16 R. Michaels please keep online and offline seperate
  • 141586 03/15/05 14:57 boeglinw n-det scaler screen
  • 141585 03/15/05 14:56 Boeglin General Hall A Tools Screen
  • 141584 03/15/05 14:45 sawatzky Computer usage comments
  • 141583 03/15/05 14:42 R. Michaels beam returns, start a run
  • 141581 03/15/05 14:22 R. Michaels adaqfs is getting full, put files on work
  • 141580 03/15/05 12:24 R. Feuerbach L-HRS S2m scaler fixed
  • 141579 03/15/05 11:25 Olivier/Brad T7 trigger rate decreased
  • 141578 03/15/05 09:42 R. Michaels deadtime studies
  • 141577 03/15/05 09:31 ran coinc neutrons TOF in C12
  • 141548 03/15/05 08:49 R. Michaels beam off, end run 2677
  • 141544 03/15/05 08:06 R. Michaels S2m Left arm rate zero
  • 141542 03/15/05 07:57 B. Craver Owl Shift Summary
  • 141533 03/15/05 05:24 ran acceptance before and after momentum change of L HRS
  • 141532 03/15/05 05:08 Ramesh Follow up of 141529
  • 141529 03/15/05 04:50 Ramesh ND P1 adc versus tdc spectrum after adc gate delayed by 12 ns
  • 141526 03/15/05 03:45 B. Craver Target screen shot (2/2)
  • 141525 03/15/05 03:44 B. Craver Target screen shot (1/2)
  • 141524 03/15/05 03:41 B. Craver General tools screen shot
  • 141523 03/15/05 03:39 B. Craver Neutron detector screen shot.
  • 141520 03/15/05 03:25 Ramesh HV of P3 left-pmt increased by 50V
  • 141515 03/15/05 01:57 Ramesh Raw ADC vs TDC of ND for Run 2665
  • 141508 03/15/05 00:02 boeglin Shift Summary
  • Monday

  • 141506 03/14/05 23:55 P. Monaghan ND and BB scaler rates
  • 141505 03/14/05 23:53 P. Monaghan Target controls screen
  • 141504 03/14/05 23:52 P. Monaghan Target loop charts
  • 141503 03/14/05 23:51 P. Monaghan N1-5 channel adc and tdc plots for runs 2655 and 2659
  • 141492 03/14/05 20:49 P. Monaghan new Compton run started
  • 141491 03/14/05 20:44 P. Monaghan N2-7L adc spectrum
  • 141490 03/14/05 20:36 P. Monaghan N1-5 adc spectra
  • 141489 03/14/05 20:27 P. Monaghan R-HRS S1 L3 adc
  • 141486 03/14/05 20:10 boeglin n-det scaler screen
  • 141485 03/14/05 20:09 Boeglin Hall A General Tool Screen
  • 141478 03/14/05 18:18 P. Monaghan Detector scaler display
  • 141477 03/14/05 18:15 P. Monaghan Harp Scan results
  • 141473 03/14/05 17:01 P. Monaghan Target: Set fan speeds to 24Hz
  • 141472 03/14/05 16:31 sawatzky LHRS momentum lowered to 3.7240 GeV/c
  • 141470 03/14/05 16:02 Saha Run Summary
  • 141469 03/14/05 14:38 P. Monaghan N2-7L base changed
  • 141468 03/14/05 14:12 Nanda Compton result
  • 141467 03/14/05 11:38 sawatzky epicRunStart: HVs are being read-out by DAQ again
  • 141466 03/14/05 11:33 sawatzky Netgear swtich by BB electronics moved to table
  • 141465 03/14/05 10:46 Olivier ND gate delayed by 12ns
  • 141464 03/14/05 10:45 Olivier N2-7L PMT is dead
  • 141463 03/14/05 10:29 saw HV EPICS control
  • 141461 03/14/05 10:02 Olivier Runs 2652 and 2653 are cosmic, not production
  • 141456 03/14/05 08:44 Ramesh Follow up of entry 141455
  • 141455 03/14/05 08:39 Ramesh Protons in BB so far
  • 141453 03/14/05 08:09 Ramesh Shift Summary
  • 141449 03/14/05 07:23 adaq BB scalers
  • 141444 03/14/05 06:12 Adaq Target status
  • 141443 03/14/05 05:58 Adaq Hall A general tools
  • 141440 03/14/05 05:31 Ramesh ROC5 failed, reset and rebooted
  • 141438 03/14/05 04:44 Ramesh Deadtime OK
  • 141436 03/14/05 04:25 Ramesh Deadtime 100 at 40uA current in C
  • 141433 03/14/05 03:42 eip scalers 40 uA carbon
  • 141432 03/14/05 03:32 eip detector replay problems
  • 141429 03/14/05 02:56 Ramesh Follow up of halog entry 141415
  • 141424 03/14/05 01:52 sawatzky New injector installed, 40 uA being delivered
  • 141423 03/14/05 01:45 Re: Commented out BB HV readout in epics startup file
  • 141422 03/14/05 01:38 sawatzky Commented out BB HV readout in epics startup file
  • 141421 03/14/05 01:29 sawatzky EPICS HV issue
  • 141419 03/14/05 01:24 sawatzky Harp scan
  • 141415 03/14/05 00:58 Ramesh Follow up of entry 141414
  • 141414 03/14/05 00:52 Ramesh Protons in BigBite from Carbon
  • 141413 03/14/05 00:23 P. Monaghan Target loop charts
  • 141412 03/14/05 00:21 P. Monaghan Target parameters screen
  • 141411 03/14/05 00:18 P. Monaghan Swing Shift Summary
  • Sunday

  • 141408 03/13/05 23:40 P. Monaghan Hall A tools screen
  • 141407 03/13/05 22:33 P. Monaghan MCC changing guns
  • 141402 03/13/05 20:32 P. Monaghan MCC injector problems
  • 141400 03/13/05 20:06 P. Monaghan ND and BB scaler display
  • 141395 03/13/05 18:16 P. Monaghan ND and BB scaler display
  • 141387 03/13/05 16:12 Boeglin Shift Summary
  • 141384 03/13/05 15:24 boeglin portserver information, n-det HV
  • 141381 03/13/05 14:33 Boeglin neg. number of T7
  • 141376 03/13/05 13:22 B. Craver Double checked HVs
  • 141375 03/13/05 13:09 Boeglin N-detector rates
  • 141373 03/13/05 12:57 sawatzky HVs part 2 of 2
  • 141372 03/13/05 12:52 sawatzky HVs part 1 of 2
  • 141371 03/13/05 10:34 Boeglin Bigbite HV problem
  • 141369 03/13/05 10:05 Adaq Hall A General Tools
  • 141368 03/13/05 10:03 B. Craver scalers
  • 141364 03/13/05 08:41 B. Craver Move to slanted carbon
  • 141363 03/13/05 08:20 Ramesh Shift Summary
  • 141362 03/13/05 08:17 eip 20uA deutron runs for analysis
  • 141359 03/13/05 07:49
  • 141356 03/13/05 06:15 Ramesh Neutrons(NA) and Protons(BB)
  • 141352 03/13/05 04:43 mazouz target parameters
  • 141349 03/13/05 03:56 Ramesh BB/NA scaler
  • 141348 03/13/05 03:45 Ramesh Hall A general tools
  • 141339 03/13/05 00:10 P. Monaghan Neutron detector channels N1-5L/R
  • 141337 03/13/05 00:03 P. Monaghan Swing Shift Summary
  • Saturday

  • 141336 03/12/05 23:56 P. Monaghan computer hacsbc2
  • 141335 03/12/05 23:47 Ramesh Neutrons in Neutron Array
  • 141332 03/12/05 23:00 saw Extracting HV settings from data file
  • 141331 03/12/05 22:34 P. Monaghan Target controls (2)
  • 141330 03/12/05 22:32 P. Monaghan Target controls screen
  • 141329 03/12/05 22:30 P. Monaghan Hall A General Tools
  • 141328 03/12/05 22:29 P. Monaghan BigBite/Neutron detector scalers
  • 141323 03/12/05 19:36 R. Michaels negative deadtimes
  • 141322 03/12/05 19:26 P. Monaghan ADC spectra for N1-5 and N1-6
  • 141318 03/12/05 17:56 sawatzky Neutron arm HV questions
  • 141317 03/12/05 17:15 Ramesh Protons in BigBite
  • 141315 03/12/05 16:29 sawatzky Change in HV for N1-5L, N1-4L
  • 141311 03/12/05 15:59 slifer Shift Summary (Day)
  • 141310 03/12/05 15:29 Boeglin neutron detector
  • 141309 03/12/05 15:24 Adaq neutron detector plane 1 pad 4
  • 141298 03/12/05 11:24 sawatzky negative deadtimes
  • 141293 03/12/05 09:21 slifer Negative deadtime
  • 141292 03/12/05 08:40 slifer Hall A tools
  • 141291 03/12/05 08:38 slifer Big Bite / ND rates
  • 141288 03/12/05 08:21 slifer target checklist
  • 141287 03/12/05 08:04 Ramesh Run Summary
  • 141284 03/12/05 06:19 mazouz LD2 target parameters
  • 141281 03/12/05 05:35 Ramesh Modified adc vs tdc plot
  • 141278 03/12/05 04:54 ran ND tdc - vs - tdc problem
  • 141277 03/12/05 04:15 Ramesh ADC vs TDC plots for ND p1
  • 141272 03/12/05 02:35 Ramesh Hall A general tools
  • 141271 03/12/05 02:33 Ramesh ND scalers at 20uA with LD2
  • 141268 03/12/05 01:15 Ramesh ND P4 adc VS tdc timing problem?
  • 141265 03/12/05 00:03 Fatiha Shift Summary
  • Friday

  • 141261 03/11/05 23:47 Kolarkar Target Screens - swing
  • 141256 03/11/05 22:23 Ramesh T8 prescaled unknowingly as 9999 for several runs
  • 141247 03/11/05 20:29 Kolarkar Detector Scaler Screen
  • 141246 03/11/05 20:28 Kolarkar Neutron Detector Single Rates Screen
  • 141245 03/11/05 20:23 Kolarkar General Tools Screen
  • 141244 03/11/05 20:22 Kolarkar Current Monitor
  • 141235 03/11/05 18:18 Fatiha Current Monitor
  • 141228 03/11/05 16:56 Fatiha FFB and Energy Lock
  • 141227 03/11/05 16:27 nanda compton
  • 141224 03/11/05 16:08 Higinbotham Shift Summary
  • 141220 03/11/05 15:37 Kelleher Big Bite scalers
  • 141219 03/11/05 15:33 Higinbotham Magnet Screen
  • 141216 03/11/05 15:19 Kelleher target screens (day shift)
  • 141214 03/11/05 14:40 Kelleher target to LD2
  • 141211 03/11/05 13:15 Kelleher 40uA BB/NA rates
  • 141210 03/11/05 13:10 Kelleher 30 uA BB/NA rates
  • 141208 03/11/05 13:04 Kelleher BB/NA rates for carbon
  • 141206 03/11/05 13:03 Kelleher target changed to carbon
  • 141204 03/11/05 12:19 Kelleher BB/NA rates
  • 141202 03/11/05 12:12 Kelleher target to LD2
  • 141201 03/11/05 09:55 Kelleher target to Optics
  • 141199 03/11/05 09:53 Higinbotham BigBite and LHRS Events
  • 141192 03/11/05 08:13 Annand Hall-A Owl Shift Summary
  • 141186 03/11/05 05:44 Thompson L-arm theta-phi assymetry
  • 141185 03/11/05 05:39 Thompson Abnormal P2-pos plots
  • 141182 03/11/05 05:19 sawatzky Power cycled BB VME crate
  • 141178 03/11/05 04:45 Annand BigBite QDC readout
  • 141177 03/11/05 04:37 sawatzky onlana.C bigbite adc plots empty
  • 141171 03/11/05 04:02 Annand BigBite & Neutron Rates 20uA
  • 141168 03/11/05 03:41 Annand BigBite & Neutron rates 15uA
  • 141165 03/11/05 03:22 Annand BigBite & Neutron Rates 10uA
  • 141162 03/11/05 03:10 Annand BigBite Rates 5uA
  • 141160 03/11/05 02:44 Neil target LD2
  • 141159 03/11/05 02:40 sawatzky Q2 crashed again
  • 141158 03/11/05 02:36 sawatzky Reboot roc5
  • 141157 03/11/05 02:35 Annand Raster++ plots
  • 141152 03/11/05 02:11 Annand Harp Scan
  • 141151 03/11/05 01:44 spiegel Hall A tech's
  • 141150 03/11/05 01:39 spiegel Hall A tech's
  • 141149 03/11/05 01:02 sawatzky Q1 still down
  • 141148 03/11/05 00:49 Ramesh Another bad harp scan
  • 141147 03/11/05 00:33 Ramesh Harp Scan
  • 141145 03/11/05 00:05 Fatiha Shift Summary
  • 141144 03/11/05 00:04 Fatiha BigBite HV tripped
  • Thursday

  • 141143 03/10/05 23:52 Fatiha Q1 and Q2 crashed
  • 141142 03/10/05 22:06 Fatiha General Tools Screen
  • 141141 03/10/05 21:47 sawatzky LRHS Q2 crashed again
  • 141139 03/10/05 21:09 sawatzky Dipole M3/M4 OverTemp Fault
  • 141138 03/10/05 20:54 Fatiha Left Q2 crashed again
  • 141137 03/10/05 20:27 Kolarkar Left Q2 back; TS reset
  • 141134 03/10/05 19:38 Fatiha Left Q2 crashed
  • 141133 03/10/05 19:35 Higinbotham Shift Summary
  • 141132 03/10/05 15:07 Kelleher target screens (day shift)
  • 141129 03/10/05 14:57 slifer HRSL to 26.1 degrees
  • 141126 03/10/05 14:50 Kelleher target to carbon
  • 141121 03/10/05 14:34 Kelleher target to LD2
  • 141118 03/10/05 14:02 Kelleher target to LH2
  • 141114 03/10/05 13:46 Kelleher 2uA BB rates
  • 141109 03/10/05 13:29 Kelleher 20uA BB rates
  • 141104 03/10/05 13:13 Kelleher 15 uA BB rates
  • 141099 03/10/05 13:06 Kelleher 10uA BB rates
  • 141097 03/10/05 13:03 slifer runs 2539-2541 are slanted carbon
  • 141093 03/10/05 12:56 Kelleher 5uA carbon BB rates
  • 141089 03/10/05 12:47 Kelleher target to slanted carbon
  • 141088 03/10/05 12:42 Higinbotham Comment on Gain Settings
  • 141085 03/10/05 12:36 Kelleher Big Bite Rates
  • 141082 03/10/05 12:32 Higinbotham dE gains too high
  • 141075 03/10/05 11:54 ran problematic channels
  • 141074 03/10/05 11:24 Higinbotham BigBite Thresholds
  • 141073 03/10/05 09:19 sawatzky LHRS Bogie control not working
  • 141071 03/10/05 07:45 Nilanga shift Summary
  • 141065 03/10/05 06:31 Nilanga Final HV settings
  • 141019 03/10/05 01:24 E.Chudakov HV settings on BigBite
  • 141017 03/10/05 00:23 E.Chudakov Shift summary
  • Wednesday

  • 140998 03/09/05 20:05 E.Chudakov Harp scan after beam adjustment
  • 140997 03/09/05 19:46 E.Chudakov Harp scan
  • 140981 03/09/05 16:49 sawatzky Dipole working again
  • 140980 03/09/05 16:45 sawatzky LHRS Dipole not settling to correct value
  • 140979 03/09/05 16:00 Higinbotham Shift Summary
  • 140977 03/09/05 15:58 sawatzky Cycled LHRS Q2/3, turned D1 back on
  • 140976 03/09/05 15:29 Higinbotham dE and E discriminator thresholds to 60 mV from 33 mV
  • 140975 03/09/05 15:29 sawatzky BB aux plane OR moved to different 'sub-module'
  • 140973 03/09/05 14:17 R. Michaels rcServer updated
  • 140955 03/09/05 08:38 sawatzky Collimator control on LHRS
  • 140954 03/09/05 08:31 slifer target checklist (day shift)
  • 140953 03/09/05 08:30 slifer target checklist (day shift)
  • 140952 03/09/05 08:29 sawatzky moveable stairs blocking LHRS motion
  • 140948 03/09/05 08:24 sawatzky reset bogie controllers
  • 140942 03/09/05 08:00 Nilanga Shift Summary
  • 140922 03/09/05 03:54 P. Monaghan Harp Scan
  • 140921 03/09/05 03:43 P. Monaghan TECH - Left arm collimator problem
  • Tuesday

  • 140919 03/08/05 19:48 Ramesh ND-scaler in cosmic
  • 140917 03/08/05 17:17 Olivier Discriminator swapped
  • 140908 03/08/05 10:50 R. Michaels ctimel fixed (for now)
  • 140905 03/08/05 10:28 atrig Trigger-Download
  • Monday

  • 140900 03/07/05 16:50 Olivier D.ctimel vs evt number
  • 140898 03/07/05 16:40 R. Michaels ctimel (Left) problems again
  • 140895 03/07/05 16:30 atrig Trigger-Download
  • 140888 03/07/05 15:02 R. Michaels L-arm TC time change
  • 140886 03/07/05 14:43 atrig Trigger-Download
  • 140885 03/07/05 14:43 atrig Trigger-Download
  • 140884 03/07/05 14:43 atrig Trigger-Download
  • 140879 03/07/05 13:10 saw V18R flakey
  • 140874 03/07/05 12:31 R. Michaels timing adjust, L-arm L1A fix
  • 140869 03/07/05 11:50 saw Pulser into SRC scalers
  • Sunday

  • 140864 03/06/05 13:15 R. Michaels DAQ recovery
  • Saturday

  • 140859 03/05/05 11:47 R. Michaels DAQ problems.
  • Friday

  • 140831 03/04/05 10:20 R. Michaels DAQ restoration
  • 140830 03/04/05 09:29 atrig Trigger-Download
  • 140828 03/04/05 09:28 atrig Trigger-Download
  • Thursday

    
    

    Wednesday

    
    

    Tuesday

  • 140797 03/01/05 16:38 P. Monaghan Change in discriminator threshold for delta-E plane
  • 140796 03/01/05 16:36 P. Monaghan Increased high voltage settings on delta-E plane